Find the flux through the top,left and right side of a cube of side 0.5 m when placed in a horizontal uniform electric field of 8.0N/C directed to the right.


1
Expert's answer
2022-02-07T08:45:02-0500

Answer

Area of top, left, right surface of cube

A=0.25m2

Electric field E=8.0N/C

Angle between all surface

"\theta=90°"

So electric flux is given in all cases

"\phi=E.A=EA cos\theta\\=8.0\times0.25\times cos0°\\=0Wb"